OpenStudy (anonymous):

4000000 /2 + 6 (32) =

Parth (parthkohli):

Can you tell me what \(4000000 \div 2\) is?

Parth (parthkohli):

Secondly, can you tell me what \(6(32)\) is?

OpenStudy (anonymous):

what are you talking about

Parth (parthkohli):

I am talking about the question that you posted above.

Parth (parthkohli):

That is what I am supposed to talk about, actually.

mathslover (mathslover):

Seems difficult wait lemme think about it .. Wait ! may i know that what you tried @bluebarbie ?

OpenStudy (anonymous):

2000192

Parth (parthkohli):

@A.Avinash_Goutham Please don't snipe questions no matter what.

OpenStudy (anonymous):

jus wondering wat wud it do after i give the answer o.o

Parth (parthkohli):

That is an irrelevant comment, bluebarbie. You are supposed to be talking about the question you had posted.

OpenStudy (anonymous):

thanks okayy

Parth (parthkohli):

Okay, now please tell me what \(4,000,000 \div 2\) is.

OpenStudy (anonymous):

Iont Know That Why I Ask For Help

mathslover (mathslover):

OK no problem : can you tell me that what is \[\frac{4}{2}\] ?

Parth (parthkohli):

\(4,000,000 \div 2\) is the same as \(\Large {4 \times 1,000,000 \over2}\)

Parth (parthkohli):

So, you may cancel the 4 and 2, and multiply the result to 1,000,000.

OpenStudy (anonymous):

4000,000/2+6*32 4000,000 divide by 2 =2000,000 and 6 multiple by 32= 192 2000,000 plus 192 total 2000,192.